An investor invested a total of $1,900 in two mutual funds. One fund earned a 5% profit while the other earned a 3% profit. If the investor's total profit was $71, how much was invested in each mutual fund?
Answer by mananth(16946)

Part I 5.00% per annum ------------- Amount invested =x
Part II 3.00% per annum ------------ Amount invested = y
1900
Interest----- 71

Part I 5.00% per annum ---x
Part II 3.00% per annum ---y
Total investment
x + 1 y= 1900 -------------1
Interest on both investments
5.00% x + 3.00% y= 71
Multiply by 100
5 x + 3 y= 7100.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -9500.00
Add this to (2)
0 x -2 y= -2400
divide by -2
y = 1200
Part I 5.00% $ 700
Part II 3.00% $ 1200

CHECK
700 --------- 5.00% ------- 35.00
1200 ------------- 3.00% ------- 36.00
Total -------------------- 71.00

m.ananth@hotmail.ca